# Calculating the EFT likelihood via saddle-point expansion

### v1

@@ Abstract
In this paper, we extend the functional approach for calculating the EFT likelihood by applying the saddle-point expansion. We demonstrate that, after suitable reformulation, the likelihood expression is consistent with the path integral required to be computed in the theory of false vacuum decay. In contrast to the saddle-point approximation, the application of the saddle-point expansion necessitates more nuanced considerations, particularly concerning the treatment of the negative eigenvalues of the second derivative of the action at the saddle point. We illustrate that a similar issue arises in the likelihood calculation, which requires approximating the original integral contour through the combination of the steepest descent contours in the field space. Gaussian distribution as a working example, we then concentrate on the calculation of the EFT likelihood and propose a general procedure for computing the likelihood via saddle-point expansion method for arbitrary partition functions. Precise computation of the likelihood will benefit Bayesian forward modeling, thereby enabling more reliable theoretical predictions.
@@ 1 Introduction
The large-scale structure (LSS) of the universe encodes valuable information about galaxy formation, structure growth, and the physics of dark matter and dark energy. To extract this information, we need to fit our theoretical predictions to the measurements from observations. Conventionally, the galaxies are regarded as the tracers of dark matter distribution, and their $n$-point correlation function is utilized to compare with data to provide constraints on cosmological parameters (see [1] for a comprehensive review).
@@
A necessary basis in applying the Bayesian method is to calculate the conditional likelihood $\mathcal{P}[\delta_{g}|\delta]$, which is the functional probability of observing the galaxy overdensity $\delta_{g}$ given the matter overdensity $\delta$ (see [19, 20] for more details). To achieve this, we need to factorize $\mathcal{P}[\delta_{g}|\delta]$ into two components, $\mathcal{P}[\delta]$ and $\mathcal{P}[\delta_{g},\delta]$, which are referred to as the matter likelihood and joint likelihood, respectively. Thanks to the recently developed path integral approach for the LSS [21, 19], both likelihoods can be naturally expressed as the functional Fourier transform of the partition functions $Z[\delta]$ and $Z[\delta_{g},\delta]$. Meanwhile, the effective field theory of the LSS (EFTofLSS) [22, 23, 21, 24, 25, 26, 27, 28, 29] postulates that the contribution of short-wavelength (small-scale) fluctuations to the long-wavelength (large-scale) universe can be considered as a sequence of non-ideal fluid effects which can be absorbed into the corresponding terms of the partition functions, such as the pressure perturbation and the viscosity. It is therefore important to establish a framework to derive the EFT likelihood (hereinafter referred to as likelihood) from a general partition function. However, for general $Z[\delta]$ and $Z[\delta_{g},\delta]$, it is usually unfeasible to evaluate $\mathcal{P}[\delta]$ and $\mathcal{P}[\delta_{g},\delta]$ through immediate calculation because of the complicated functional integral formula. In Ref. [19], a functional approach for computing the likelihood through the saddle-point approximation is proposed, yielding results that are nearly consistent with the traditional EFT estimate, except for additional higher-order terms [20]. Furthermore, these higher-order terms are identified as precise corrections to the likelihood expression.
@@
The present paper aims to calculate the likelihood more precisely. Since the saddle-point contribution in the exponent provides a good approximation to the full likelihood, it is natural to extend this approach to higher orders around the saddle points, i.e. the saddle-point expansion. This semiclassical method was first proposed in quantum mechanics and quantum field theory to calculate the decay rates and leads to reliable results that are consistent with the quantum-mechanical estimates of tunneling rates [30, 31]. More recently, the saddle-point expansion method has also been applied to Yang-Mills theory, leading to the CP conservation in the strong interactions [32, 33]. In this paper, we expect to apply this method within the path-integral formulation for LSS to achieve a more accurate prediction of the conditional likelihood.
@@
In the process of utilizing the saddle-point expansion, there is a notorious problem corresponding to the negative eigenvalues of the second derivative of the classical action at the saddle point. If we substitute the saddle-point solution into the integral, this mode will change the sign on the exponent of the path integral, thus making the integral ill-defined. The resolution to the problem is to apply the Picard-Lefschetz theory [34, 35], deforming the integral contour into the combination of the steepest descent contours which connect different saddle points, and terminate at the convergent regions of the integral. In this way, the imaginary part of different “sub-contours” (often referred to as the Lefschetz thimbles) will cancel each other out to give the real result [36]. In this paper, we will show that we also encounter negative modes, and even complex modes in the calculation of the likelihood, and we can make use of the steepest descent contour method to guarantee the result real. This is the kernel of our work because the result has the physical meaning of probability.
@@
One of our main results is that the saddle-point expansion method used in the calculation of decay rates can be perfectly applied to the calculation of the likelihood. What we need to do is just rewriting the likelihood expression into the following form
@@
where $\boldsymbol{\phi}_{g}$ is the field that absorbs all the integral parameters and $S_{g}[\boldsymbol{\phi}_{g}]$ is the “modified action”. Then we can use the same arguments as in quantum mechanics and quantum field theory to solve the expression of the likelihood analytically. At the same time, a problem that is not the same as the situation in quantum field theory is that the action is often not real because its expression often contains the imaginary unit $i$. However, in Sec. 4 we will prove that although $S_{g}[\boldsymbol{\phi}_{g}]$ and the eigenvalues of $S_{g}^{\prime\prime}[\bar{\boldsymbol{\phi}}_{g}]$ may be complex, we can still guarantee the integral result given by the sum of different steepest descent contours is real.
Also, we will illustrate how to apply the saddle-point expansion method to compute the likelihood for a given $S_{g}[\boldsymbol{\phi}_{g}]$ by a concrete example, i.e. the EFT likelihood with Gaussian distribution. We will provide a detailed introduction to how to use the gradient flow equation to find the steepest descent contours in the field space. We emphasize here although we have only studied a specific model, our discussion of this method and the treatment of the negative (complex) modes do not depend on the expressions of the action. Our approach is general and can be applied to more complicated situations.
The organization of this paper is as follows. In Sec. 2, we review the saddle-point expansion method in quantum mechanics and quantum field theory to calculate the decay rates. The analogy between quantum field theory and LSS path-integral case will be drawn in Sec. 3. Sec. 4 contains an example of applying this method to calculate the likelihood. Then we will draw our conclusion and discuss future directions in Sec. 5. To make our work complete, we review how to calculate the contribution of the saddle points to the likelihood in Appendix A, and Appendix B contains some discussions on how to solve the eigenvalues of $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$.
Throughout this paper, the notion and conventions used in the path-integral approach of the LSS mainly come from [19].
@@ 2 Saddle-point expansion in quantum mechanics and quantum field theory
In this section, we will review the application of saddle-point expansion in quantum mechanics and quantum field theory, inspired by the calculation of decay rates [30, 31, 36, 35]. We will see that the existence of the negative modes of the second derivative of the action at the saddle points compels us to complexify the integral path and deform the integral contour. We need to approximate the integral contour by the sum of the steepest descent contours connecting different saddle points, which is referred to as the combination of different Lefschetz thimbles in the Picard-Lefschetz theory [34, 37]. Our review contains only the most central parts of the saddle-point expansion method, while more detailed and rigorous reviews can be found in [36, 35].
@@ 2.1 Quantum-mechanical case
where $\bar{x}$ is the stationary point of the action, satisfying $(\delta S[{x}]/\delta x)|_{x=\bar{x}}=0$, equivalent to the classical equation of motion
@@
where the prime denotes differentiation with respect to $x$. The classical part will dominate the path integral (if we recover $\hbar$ in the theory, then the coordinate expansion should be $x=\bar{x}\,+\,\hbar^{1/2}\tilde{x}$, see for example in [38]), thus we can expand the action around the stationary point $x=\bar{x}$ and the path integral can be formulated as
$\int\mathcal{D}x(\tau)\,e^{-S_{E}[x(\tau)]}\,$ $\approx\int\mathcal{D}\tilde{x}(\tau)\,e^{-S_{E}[\bar{x}]-\frac{1}{2}S^{\prime\prime}_{E}[\bar{x}]\tilde{x}^{2}},$ (2.4)
@@
where on the right-hand side (RHS) the linear term vanishes because $\bar{x}$ is the solution of the equation of motion (given a general action, there is usually more than one solution $x_{n}$, we need to sum over all the contributions). The issue emerges when we proceed to the subsequent step: to evaluate the RHS of Eq. (2.4), it is necessary to determine the eigenvalues of $S^{\prime\prime}[\bar{x}]$. However for the single bounce solution [^1] In the false vacuum decay theory, the bounce solution refers to a trajectory that starts with zero velocity, bounces out the potential well for $n$ times, and eventually returns to the origin again, as illustrated in [30]. Since the bounce solution exists only in Euclidean space rather than Minkowski spacetime, it is also sometimes referred to as an instanton. of Eq. (2.3) there is always a negative mode, making the path-integral ill-defined [31]. There is also a zero mode proportion to $dx/d\tau$. To handle the zero mode, we can convert it into a collective coordinate [39, 36, 40]. However, the treatment of the negative mode involves several subtle aspects, which will be discussed in greater detail below.
@@
The problem requires the application of the steepest descent contour method. The main strategy consists of the following steps: (1) We complexify the action and identify all the complex saddle points, denoted by $s_{1},s_{2}\cdots,s_{n}$. (2) For each $s_{i}$ in the complex plane, we determine a corresponding steepest descent contour $C_{i}$  [^2] We emphasize that the choice of the steepest descent contour is not unique. In fact, all contours that render the integral finite are called integration cycles. These cycles and their equivalence relations form a homology group. What is required is the contour that is homologous to the original one, as illustrated in [35]. (also known as the Lefschetz thimble [34]). Concretely, each $C_{i}$ is defined simply by moving away from $s_{i}$ in the direction that increases the real part of $S$ as quickly as possible. (3) Following the above spirit, the Lefshcetz thimble corresponding to each $s_{i}$ either terminates at the convergent regions of the integral or another saddle point. Therefore, each thimble provides a convergent path integral along the complex integration contour in field space. (4) With these thimbles in field space, we can approximate the integral contour as the combination of different thimbles fulfilling the conditions, each thimble corresponds to
@@
Figure 2: The saddle points of Eq. (2.7) and a choice of the steepest descent contour in the complexified $z$-plane. Here, the yellow parentheses represent the convergent regions of the integral, while the different colored paths correspond to the integral contours passing through different saddle points.
To illustrate the process, consider a specific example. Suppose the action takes the form
@@
and we want to integrate it as in Eq. (2.1) along the real line. This function has three (complex) saddle points at $z=-1,0,1$, labeled as $s_{-1},s_{0},s_{1}$. The method then instructs us to find the Lefschetz thimbles associated with each saddle point. Fig. 2 shows the saddle points in the complex plane and a choice of the steepest descent contour for each saddle point, in which the parentheses represent the convergent regions of the path integral. It can be seen that the real line can be approximated by a combination of three Lefschetz thimbles, i.e. $C\approx C_{-1}+C_{0}+C_{1}$. Thus the path integral can be expressed as
$\mathcal{Z}$ $=\mathcal{J}_{-1}+\mathcal{J}_{0}+\mathcal{J}_{1}$
$\sim\mathcal{I}_{-1}+\mathcal{I}_{0}+\mathcal{I}_{1}\,.$ (2.8)
@@
$\mathcal{I}_{-1}=\sqrt{\pi}\,{\rm exp}(\frac{1}{4})\,,\,\,\,\mathcal{I}_{0}=\sqrt{-2\pi}\,,\,\,\,\mathcal{I}_{-1}=\sqrt{\pi}\,{\rm exp}(\frac{1}{4})\,.$ (2.9)
A noteworthy point is that the sum of these three components does not appear to be real, as the contribution of $\mathcal{I}_{0}$ is imaginary. However, this arises because we have used $\mathcal{I}_{i}$ to approximate the exact integral $\mathcal{J}_{i}$, rather than the result of applying the saddle-point expansion. If we rigorously calculate the integral through $\mathcal{J}_{i}$, although the process for $\mathcal{J}_{0}$ there will be an imaginary part, the contribution from the other two components $\mathcal{J}_{-1}$ and $\mathcal{J}_{1}$ along the imaginary axis cancels them exactly, guaranteeing the full path integral real. The most general method for solving the steepest descent contour will be illustrated in Sec. 4.

### v2

@@ Abstract
In this paper, we extend the functional approach for calculating the EFT likelihood by applying the saddle-point expansion. We demonstrate that, after suitable reformulation, the likelihood expression is consistent with the path integral required to be computed in the theory of false vacuum decay. In contrast to the saddle-point approximation, the application of the saddle-point expansion necessitates more nuanced considerations, particularly concerning the treatment of the negative eigenvalues of the second derivative of the action at the saddle point. We illustrate that a similar issue arises in the likelihood calculation, which requires approximating the original integral contour through the combination of the steepest descent contours in the field space. As a concrete example, we focus on calculating the EFT likelihood under a Gaussian distribution and propose a general procedure for computing the likelihood using the saddle-point expansion method for arbitrary partition functions. Precise computation of the likelihood will benefit Bayesian forward modeling, thereby enabling more reliable theoretical predictions.
@@ 1 Introduction
The large-scale structure (LSS) of the universe encodes valuable information about galaxy formation, structure growth, and the physics of dark matter and dark energy. To extract this information, we need to fit our theoretical predictions to the measurements from observations. Conventionally, the galaxies are regarded as the tracers of the dark matter distribution, and their $n$-point correlation function is used to compare with data to provide constraints on cosmological parameters (see Ref. [1] for a comprehensive review).
@@
A necessary basis in applying the Bayesian method is to calculate the conditional likelihood $\mathcal{P}[\delta_{g}|\delta]$, which is the functional probability of observing the galaxy overdensity $\delta_{g}$ given the matter overdensity $\delta$ (see Refs. [19, 20] for more details). To achieve this, we need to factorize $\mathcal{P}[\delta_{g}|\delta]$ into two components, $\mathcal{P}[\delta]$ and $\mathcal{P}[\delta_{g},\delta]$, which are referred to as the matter likelihood and joint likelihood, respectively. Thanks to the recently developed path integral approach to the LSS [21, 19], both likelihoods can be naturally expressed as the functional Fourier transform of the partition functions $Z[\delta]$ and $Z[\delta_{g},\delta]$. Meanwhile, the Effective Field Theory of LSS (EFTofLSS) [22, 23, 21, 24, 25, 26, 27, 28, 29] postulates that the contribution of short-wavelength (small-scale) fluctuations to the long-wavelength (large-scale) universe can be considered as a sequence of non-ideal fluid effects which can be absorbed into the corresponding terms of the partition functions, such as the pressure perturbation and the viscosity. It is therefore important to establish a framework to derive the EFT likelihood (hereinafter referred to as likelihood) from a general partition function. However, given general $Z[\delta]$ and $Z[\delta_{g},\delta]$, it is usually unfeasible to evaluate $\mathcal{P}[\delta]$ and $\mathcal{P}[\delta_{g},\delta]$ by direct calculation because of the complicated functional integral formula. In [19], a functional approach for computing the conditional likelihood through the saddle-point approximation is proposed, yielding results that are nearly consistent with the traditional EFT estimate, except for additional higher-order terms [20]. Furthermore, these higher-order terms are identified as precise corrections to the conditional likelihood expression.
@@
The present paper aims to calculate the likelihoods more precisely. Since the saddle-point contribution in the exponent provides a good approximation to the full likelihood, it is natural to extend this approach to higher orders around the saddle points, i.e., via the saddle-point expansion. This semiclassical method was first proposed in quantum mechanics and quantum field theory to calculate the decay rates and leads to reliable results that are consistent with the quantum-mechanical estimates of tunneling rates [30, 31]. More recently, the saddle-point expansion method has also been applied to Yang-Mills theory, leading to the CP conservation in the strong interactions [32, 33]. In this paper, we expect to apply this method within the path-integral formulation of LSS to achieve a more accurate prediction of the conditional likelihood.
@@
In the process of applying the saddle-point expansion, there is a notorious problem corresponding to the negative eigenvalues of the second derivative of the classical action at the saddle point. When we substitute the saddle-point solution into the integral, this mode will change the sign on the exponent of the path integral, thus making the integral ill-defined. The resolution to the problem, is to apply the Picard-Lefschetz theory [34, 35], deforming the integral contour into the combination of the steepest descent contours which connect different saddle points, and terminate at the convergent regions of the integral. In this way, the imaginary components of different “sub-contours” (often referred to as the Lefschetz thimbles) will cancel each other out to give the real result [36]. In this paper, we will show that we also encounter negative modes, and even complex modes in the calculation of the likelihoods, and we can make use of the steepest descent contour method to guarantee the result real. This is the kernel of our work because the result has the physical meaning of probability.
@@
One of our main results is that the saddle-point expansion method used in the calculation of decay rates can be perfectly applied to the calculation of the likelihoods. What we need to do is simply reformulate the likelihoods into the appropriate form. Take the case of the joint likelihood as an example, we can write
@@
where $\boldsymbol{\phi}_{g}$ is the field that absorbs all the integral parameters and $S_{g}[\boldsymbol{\phi}_{g}]$ is the “modified action”. Then we can use the same arguments as in quantum mechanics and quantum field theory to solve the expression for the likelihood analytically. At the same time, a problem that is not the same as the situation in quantum field theory is that the action is often not real because its expression often contains the imaginary unit $i$. However, in Sec. 4 we will prove that although $S_{g}[\boldsymbol{\phi}_{g}]$ and the eigenvalues of $S_{g}^{\prime\prime}[\bar{\boldsymbol{\phi}}_{g}]$ may be complex, we can still guarantee the integral result given by the sum of different steepest descent contours is real.
Also, we will illustrate how to apply the saddle-point expansion method to compute the likelihood for a given $S_{g}[\boldsymbol{\phi}_{g}]$ by a concrete example, i.e. the EFT likelihood with a Gaussian distribution. We will provide a detailed introduction on how to use the gradient flow equation to find the steepest descent contours in the field space. We emphasize here that, although we have only studied a specific model, our discussion of this method and the treatment of the negative (complex) modes do not depend on the form of the action. Our approach is general, and can be applied to more complicated situations.
The organization of this paper is as follows. In Sec. 2, we review the saddle-point expansion method in quantum mechanics and quantum field theory to calculate the decay rates. The analogy between quantum field theory and LSS path-integral case will be drawn in Sec. 3. Sec. 4 contains an example of applying this method to calculate the conditional likelihood. Then we will draw our conclusion and discuss future directions in Sec. 5. To complete our work, we provide a direct calculation illustrating the cancellation of the imaginary contributions arising from different steepest descent contours in Appendix A, and we also review how to calculate the contribution of the saddle points to the likelihood in Appendix B. We present detailed discussions on how to solve the eigenvalues of $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$ in Appendix C.
Throughout this paper, the notations and conventions used in the path-integral approach to the LSS mainly come from [19].
@@ 2 Saddle-point expansion in quantum mechanics and quantum field theory
In this section, we will review the application of the saddle-point expansion in quantum mechanics and quantum field theory, inspired by the calculation of decay rates [30, 31, 36, 35]. We will see that the existence of the negative modes of the second derivative of the action at the saddle points compels us to complexify the integral path and deform the integral contour. We need to approximate the integral contour by the sum of the steepest descent contours connecting different saddle points, which is referred to as the combination of different Lefschetz thimbles in the Picard-Lefschetz theory [34, 37]. Our review contains only the most central parts of the saddle-point expansion method, while more detailed and rigorous reviews can be found in [36, 35].
@@ 2.1 Quantum-mechanical case
where $\bar{x}$ is the stationary point of the action, satisfying $(\delta S[{x}]/\delta x)|_{x=\bar{x}}=0$, equivalent to the classical equation of motion (EOM)
@@
where the prime denotes differentiation with respect to $x$. It is evident that the EOM in Euclidean space, compared to its counterpart in Minkowski spacetime, is equivalent to a sign flip of the potential term. Consequently, in the context of general false vacuum decay theory, this EOM Eq. (2.3) typically admits three solutions: (1) The trivial false-vacuum static solution $x_{F}(\tau)\equiv x_{+}$. (2) The bounces solution $x_{B_{n}}(\tau)$ [^1] In quantum field theory, since the bounces solution exists only in Euclidean space rather than Minkowski spacetime, it is also known as the instanton., which refers to a trajectory that starts with zero velocity, bounces out of the potential well $n$ times, and eventually returns to the origin again, as illustrated in [30]. (3) The last one is the shot solution $x_{S}(\tau)$, described in [38, 36], which quickly reaches the other side of the potential well at $V(x)=0$, where it stays for the majority of the time, before ultimately returning to the origin. These three solutions are vividly illustrated in Fig. 5 of [36]. Notably, the bounces solution does not correspond to a local minimum of the action but rather to a saddle point. To emphasize the essence of the saddle-point expansion, we omit the detailed summation over the multi-bounces and subsequently denote it succinctly as $B$ or bounce (the details for the summation process can be found in, e.g., [35]).
@@
The classical part will dominate the path integral (if we recover $\hbar$ in this theory, then the coordinate decomposition should be $x=\bar{x}\,+\,\hbar^{1/2}\tilde{x}$, see for example in [39]), thus we can expand the action around all the stationary points $x=\bar{x}_{n}$ ($n=FV$, $B$, and $S$). Then the path integral can be formulated as
$\int\mathcal{D}x(\tau)\,e^{-S_{E}[x(\tau)]}\,\approx\sum_{n=FV,B,S}\int\mathcal{D}\tilde{x}_{n}(\tau)\,e^{-S_{E}[\bar{x}_{n}]-\frac{1}{2}S^{\prime\prime}_{E}[\bar{x}_{n}]\tilde{x}_{n}^{2}},$ (2.4)
@@
where on the right-hand side (RHS) the linear term vanishes because $\bar{x}_{n}$ are the solutions of the EOM. The issue emerges when we proceed to the subsequent step: to evaluate the RHS of Eq. (2.4), it is necessary to determine the eigenvalues of $S_{E}^{\prime\prime}[\bar{x}]$. However, for the single bounce solution of Eq. (2.3) there is always a negative mode, making the path-integral ill-defined [31]. There is also a zero mode proportion to $dx/d\tau$. To handle the zero mode, we can convert it into a collective coordinate [40, 36, 41]. However, the treatment of the negative mode involves several subtle aspects, which will be discussed in greater detail below.
@@
In fact, the resolution of this problem requires the application of the steepest descent contour method. The main strategy consists of the following steps: (1) We complexify the action and identify all the complex saddle points, denoted by $s_{1},s_{2}\cdots,s_{n}$. (2) For each $s_{i}$ in the complex plane, we determine a corresponding steepest descent contour $C_{i}$ [^2] We emphasize that the choice of the steepest descent contour is not unique. In fact, all contours that render the integral finite are called integration cycles. These cycles and their equivalence relations form a homology group. What is required is the contour that is homologous to the original one, as illustrated in [35]. (also known as the Lefschetz thimble [34]). Concretely, each $C_{i}$ is defined simply by moving away from $s_{i}$ in the direction that increases the real part of $S$ as quickly as possible. (3) Following the above spirit, the Lefshcetz thimble corresponding to each $s_{i}$ either terminates at the convergent regions of the integral or another saddle point. Therefore, each thimble provides a convergent path integral along the complex integration contour in field space. (4) With these thimbles in field space, we can approximate the integral contour as the combination of different thimbles fulfilling the conditions, each thimble corresponds to
@@
Figure 2: The saddle points of Eq. (2.7) and a choice of the steepest descent contours in the complexified $z$-plane. Here, the yellow parentheses represent the convergent regions of the integral, while the different colored paths correspond to the integral contours passing through different saddle points.
To illustrate the process clearly, consider a specific example. Suppose the action takes the form
@@
and we want to integrate it as in Eq. (2.1) along the real line. This action has three (complex) saddle points at $z=-1,0,1$, labeled as $s_{-1},s_{0},s_{1}$. The method then instructs us to find the Lefschetz thimbles associated with each saddle point. Fig. 2 shows the saddle points in the complex plane and a choice of the steepest descent contour for each saddle point, in which the parentheses represent the convergent regions of the path integral. Simultaneously, the dashed line in the figure represents the original integral contour $C=\mathbb{R}$. From the figure, we can observe that $C_{-1}$ is the path that originates from the left extremity of the real line, passes through the saddle point $s_{-1}$, reaches the origin, and then proceeds downward along the imaginary axis toward ${\rm Im}z\rightarrow-\infty$, satisfying the requirement of the steepest descent contour. The other two paths can be interpreted in the same manner. We hereby briefly explain why the combination of the three contours is homologous to the original one: since the entire complex plane is effectively spherical, the endpoint of the contour $C_{-1}$ at ${\rm Im}z\rightarrow-\infty$ is connected to the starting point of $C_{0}$ at ${\rm Im}z\rightarrow\infty$. A similar connection also occurs between the paths $C_{0}$ and $C_{1}$. Thus, the sum of these three components, $C_{-1}+C_{0}+C_{1}$, forms a trajectory that originates from ${\rm Re}z\rightarrow-\infty$, traverses through the three saddle points, and ultimately terminates at ${\rm Re}z\rightarrow\infty$. It is evident that this path can be obtained through a continuous deformation of the original path $C=\mathbb{R}$, implying that they belong to the same equivalent class within the homology group of the manifold. Therefore, as proposed in the Picard-Lecschetz theory, the integral along the real line can be approximated by the combination of the three Lefschetz thimbles, i.e., $\int_{C}\approx\int_{C_{-1}}+\int_{C_{0}}+\int_{C_{1}}$, and the path integral can be formulated as
$\mathcal{Z}=\mathcal{J}_{-1}+\mathcal{J}_{0}+\mathcal{J}_{1}\sim\mathcal{I}_{-1}+\mathcal{I}_{0}+\mathcal{I}_{1}\,.$ (2.8)
@@
$\mathcal{I}_{-1}=\sqrt{\pi}\,{\rm exp}\left(\frac{1}{4}\right)\,,\,\,\,\mathcal{I}_{0}=\sqrt{-2\pi}\,,\,\,\,\mathcal{I}_{-1}=\sqrt{\pi}\,{\rm exp}\left(\frac{1}{4}\right)\,.$ (2.9)
A noteworthy point is that the sum of these three components does not appear to be real, as the contribution of $\mathcal{I}_{0}$ is imaginary. However, this arises because we have used $\mathcal{I}_{i}$ to approximate the exact integral $\mathcal{J}_{i}$, rather than the result of applying the saddle-point expansion. If we rigorously calculate the integral through $\mathcal{J}_{i}$, although the process for $\mathcal{J}_{-1}$ will generate an imaginary part, the contribution from the other component $\mathcal{J}_{1}$ along the imaginary axis cancels it exactly, guaranteeing that the full path integral remains real. We provide a detailed discussion of this cancellation process in Appendix A, where we also calculate the sum of different integrals exactly along all Lefschetz thimbles, which is
$\mathcal{J}_{-1}+\mathcal{J}_{0}+\mathcal{J}_{1}\approx\exp\left(\frac{1}{4}\right)\left(\sqrt{\pi}+0.75\right)+\sqrt{2\pi}\,,$ (2.10)
Subsequently, whenever we apply the saddle-point expansion method to compute integrals, we will always use the exact integral $\mathcal{J}_{i}$ rather than $\mathcal{I}_{i}$. In Sec. 4, we introduce a general approach that employs the gradient flow equation to find the steepest descent contours.